calculate the PH of 20 ml 0.25M CH3COOH and 50ml 0.1M NaOH solution

Grade:

1 Answers

Arun
4 years ago
  • The answer is simple.
  • pH= pKa+ log (salt/acid).
  • In the mixture, No.of m.moles of acetic acid=20 x0.25= 5 m moles.
  • The no.of m moles of NaOH =50x0.1= 5 m.moles.
  • So log (salt/acid)= (log1/0) =0
  • So pH= pKa= 4.74